10:30A – 8:ooP
- Stop O/N confocal by 11A.
- Confocal data check: Late positions start too low. Not sure why — maybe middle position to close to edge changed microscope positions?
- Collect virgin Espl[D]/ sim[D].
- Flip fly stocks (3/4ths complete)
- Flip fly cages, 12 noon
- some larvae hatched on both Psc and Pc plates (not 100% penetrant knockdown at least. Now to see if we get any phenotypes at all…)
- 5PM, collect virgin
- Fix Pc and Psc knockdown embryos
- Fraction of embryos saved at -20C
- rehydrate embryos, block, O/N incubation in m anti-AbdB
- Fly cages continuing to lay excellently.
DNA FISH probe making
- spin down and miniprep colonies. Plasmid Concentrations look good (250-350ng/uL)
- Check O/N probe synthesis run. No short probes detected. General band in 647: either loading dye or un-encorporated nucleotides.
- Run test reaction with new 5kb plasmid DNA.
- Testing reaction on gel. — Typhoon not talking to computer. But looks okay on Sapphire gel box.
- Out of dGAC nucleotide mix for labeling reaction. Got old frozen 100 mG GTPs from Sara. Making new dilute nucleotide and nucleotide mix solutions.
- Set up O/N DFISH probe synthesis.
- Fix S2 cells on coverslip. Prepping for DNA FISH with small volume.
- DNA FISH labeling at 37 O/N with 25uL of probe between coverslip and slide. (we’ll see if this can be removed and washed without detaching the cells…)
RNA FISH in culture cells, troubleshooting
- Fix cells in 8 well slide
- Treat in 100% MeOH at room temp 6 hrs, then O/N at 4C. (permiabilize/dissolve membranes to allow better probe penetration).
- Confocal 41 positions of MP10 sna/y.
- top/bottom looking good, still on position 2 (good 10 sections of clearance above and below out of 60 section .31 um steps).
- not looking on track to finish all 41 before 1PM tomorrow, increased scan speed to 7, dot resolution still looks pretty sharp on the 780.
- now going at about 3 per hour, positioning still looks okay.